Find the square root of 7. Does it repeat? does it end? Is it a irrational or rational number?
1
Expert's answer
2020-09-01T14:34:12-0400

If we use a calculator, we'll obtain "\sqrt{7} \approx 2.64575131106459."


Let us first determine if the square root of 7 is rational or irrational. Let us assume it is rational, so

"\sqrt{7} = \dfrac{m}{n} \Rightarrow 7 = \dfrac{m^2}{n^2} \Rightarrow m^2 = 7n^2." m and n are integers, so "m^2" is divisible by 7, so m is divisible by 7. Therefore,

"(7k)^2 = 7n^2 \Rightarrow 7k^2 = n^2." So n should be divisible by 7, and so on. We will obtain the same equality, but m and n can't decrease infinitely. Therefore, our assumption is incorrect, so "\sqrt{7}" is irrational.


If the fraction ends, we can write it in the form "a_0.a_1a_2a_3\ldots a_n" . But we can transform it into a rational number

"a_0.a_1a_2a_3\ldots a_n = \dfrac{a_0a_1a_2a_3\ldots a_n}{10^n}" . However, we proved that "\sqrt{7}" is irrational, so it can't be represented as a finite fraction.


If the fraction repeats, we call it periodical. Let it have a period of n numbers: "\sqrt{7} = X = a_0.(a_1\ldots a_n)" . Let us prove it is a rational number.

"10^nX - X = a_0a_1\ldots a_n.(a_1\ldots a_n) - a_0.(a_1\ldots a_n) = a_0a_1\ldots a_n - a_0."

"X(10^n-1) = a_0a_1\ldots a_n - a_0 \Longrightarrow X = \dfrac{a_0a_1\ldots a_n - a_0 }{10^n-1}."

So our X is rational. However, we proved that "\sqrt{7}" is irrational, so it can't be represented as a periodical fraction.